The odds of rain in June are 13:7. If rain on successive days is independent
a.Find the probability that it will rain


b. Find the probability that it will rain in 3 days in a row


Answer by stanbon(75887)

The odds of rain in June are 13:7. If rain on successive days is independent
a.Find the probability that it will rain
P(rain) = 13/(13+7) = 13/20
-----------------------------------
b. Find the probability that it will rain in 3 days in a row
P(rain for 3) = (13/20)^3 = 0.2746
======================
Cheers,
Stan H.
==============